OpenStudy (anonymous):

Determine this product. (-3.3) x 6

OpenStudy (alexandervonhumboldt2):

-3.3*6=-(3.3*6)=solve

OpenStudy (anonymous):

..uh

OpenStudy (alexandervonhumboldt2):

you can't multiply 3.3*6?

OpenStudy (anonymous):

i did

OpenStudy (alexandervonhumboldt2):

what did you get?

OpenStudy (anonymous):

19.8

OpenStudy (alexandervonhumboldt2):

correct but the initial problem was -3.3*6=-(3.3*6)=-19.8